Foggs Tall Ship Cruises is a single-price monopoly. The table gives Foggs demand schedule nd marginal naveue Price (dollars per Quantity cruise (cruises per hour) 220 200 180 160 140 120 Draw points that show the quantity demanded 1) when the price of a cruise is S220. Label it 1 2) when the price of a cruise is S120. Label it 2 Draw the demand curve through the points. Label it D. Draw points on the marginal revenue curve 1) when the quantity increases from 0 to 1. Label it 3. 2) when the quantity increases from 4 to 5. Label it 4. Draw the marginal revenue curve. Extend it to the y-axis and label it MR Quanity (onuises per hour) >>>Draw only the objects specified in the question.
Answer #1
Price Quantity TR MR
220 0 0 -
200 1 200 200 - 0 = 200
180 2 360 360 - 200 = 160
160 3 480 480 - 360 = 120
140 4 560 560 - 480 = 80
120 5 600 600 - 560 = 40

nd marginal naveue Quanity (onuises per hour)